OpenStudy (anonymous):

what is the limit as x approaches 0 of sin^2x/x

OpenStudy (anonymous):

what method do you have at your disposal?

OpenStudy (anonymous):

graph, then prove algebraically, I could subsitute in 0.. but that just makes it a mess

OpenStudy (anonymous):

it's sin^(2)x btw

OpenStudy (anonymous):

actually we do not need much \[\frac{\sin^2(x)}{x}=\sin(x)\times \frac{\sin(x)}{x}\]

OpenStudy (anonymous):

as \(x\to 0\) we get \(\sin(x)\to\sin(0)=0\) since sine is continuous we can just plug in the number

OpenStudy (anonymous):

and a well known limit is that \[\lim_{x\to 0}\frac{\sin(x)}{x}=1\] so we get \(0\times 1=0\)

OpenStudy (anonymous):

here is a nice picture, you can see it goes right through \((0,0)\) http://www.wolframalpha.com/input/?i=sin^2%28x%29%2Fx

OpenStudy (anonymous):

I'm confused about as x approaches 0 we get sinx to sin0, that confuses me.

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

most all function you know are continuous on their domains. that is to say, if you want to compute a limit of \(f\) as say \(x\to a\) you would simply compute \(f(a)\)

OpenStudy (anonymous):

if \(a\) is not in the domain, then you have to do something else but for example if i want \[\lim_{x\to 2}\frac{x^2+x}{x+1}\] i would compute \[\frac{2^2+2}{2+1}\] and be done with it

OpenStudy (anonymous):

now you have \[\sin(x)\times\frac{\sin(x)}{x}\]and you want the limit as x goes to zero 0 is not in the domain of this thing, because you have \(x\) in the denominator but one thing you probably have seen is that \[\lim_{x\to 0}\frac{\sin(x)}{x}=1\]

OpenStudy (anonymous):

and by what i wrote above, since \(\sin(x)\) is continuous everywhere, we know \(\lim_{x\to a}\sin(x)=\sin(a)\) so in particular \(\lim_{x\to 0}\sin(x)=\sin(0)=0\)

OpenStudy (anonymous):

in other words the limit of the function is equal to the value of the function now armed with the knowledge that \(\lim_{x\to 0}\frac{\sin(x)}{x}=1\) and also knowing that "the limit of a product is the product of the limits" we get \[\lim_{x\to 0}\sin(x)\times \frac{\sin(x)}{x}=0\times 1=0\]